In a certain population, it is claimed that the mean number of years of education is 13.2, while the standard deviation is 2.95 years. A random sample of 60 people is drawn from this population, and the sample mean is 13.87 years. What are the hypotheses and test statistic to be used in this case?


1
Expert's answer
2022-06-08T14:24:17-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=13.2"

"H_1:\mu\not=13.2"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{13.87-13.2}{2.95/\sqrt{60}}=1.76"

Since it is observed that "|z|=1.76<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>1.76)= 0.078408," and since "p=0.078408>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 13.2, at the "\alpha = 0.05" significance level.